A closed box is filled with dry ice at a temperature of -90.4 oC, while the outside temperature is 25.8 oC. The box is cubical, measuring 0.335 m on a side, and the thickness of the walls is 4.13 × 10-2 m. In one day, 3.58 × 106 J of heat is conducted through the six walls. Find the thermal conductivity of the material from which the box is made.
Answer #1

siven L = L= A = thickness of side 4.13 x 10²m = Area of side 0.0413m (0.335m) 2 0.112225 m² AT = Ti-T2 = 25.8 +90 4 =116.2 K